OpenStudy (anonymous):

find the equation for a hyperbola centered at (0,0),with foci at (-square root 34,0) and (square root 34,0) and vertices at (3,0) and (-3,0)

OpenStudy (anonymous):

THe hyperbola is horizontally oriented. c^2 = a^2 + b^2 34 = 9 + b^2 b^2 = 25 b = 5 x^2/9 - y^2/25 = 1